Question : 1 A. 20 nm- B. 44mm = C. 5mm = D. 15 cm² = E. 2 cm =
Answer #1

A. 20 nm = 20×10-6 mm

B. 44mm = 44 × 10-6 Km

C. 5mm = 0.5 Cm

D. 15 cm2 = 15 × 10-4  m2

E. 2 cm3  = 2× 10 -6 m3